a plant from species A 2n=6 fertilizes a plant from a species G 2n=8. if a complete non disjunction occurs during meiosis of A while G produces normal gemeted how many chromosomes could the offsprings have
Answer #1

Species A 2n=6 (Non-disjunction results in double the expected number of chromosomes in a cell.
e.g. after meiosis the expected chromosome number in the gamete cell is n , but if non-dysjunction occurs during meiosis then the gamete will end up with 2n that is twice the expected number of chromosomes).

Therefore the expected number of chromosome due to non disjunction in (species A) 2n = 6

Species G 2n=8

As it is normal so due to meiosis the expected number of chromosome n in (species G) n = 8/2 = 4

Therefore,

OFFSPRINGS WILL HAVE NUMBER OF CHROMOSOME 3n= 4+6= 10

3n= 4+6= 10 (3n = expected number of chromosome in species G + expected number of chromosome in species A)

If non-disjunction produces a diploid (2n) gamete and it then fuses with a normal haploid (n)
gamete, then a triploid (3n) organism would be formed (but it would be infertile as each
chromosome would not have a homologous pair at meiosis).